### v2

# A characterization of $(\mu,\nu)$-dichotomies via admissibility

Abstract.
We present a characterization of $(\mu,\nu)$-dichotomies in terms of the admissibility of certain pairs of weighted spaces for nonautonomous discrete time dynamics acting on Banach spaces. Our general framework enables us to treat various settings in which no similar result has been previously obtained as well as to recover and refine several known results. We emphasize that our results hold without any bounded growth assumption and the statements make no use of Lyapunov norms. Moreover, as a consequence of our characterization, we study the robustness of $(\mu,\nu)$-dichotomies, i.e. we show that this notion persists under small but very general linear perturbations.
Key words and phrases:
discrete nonautonomous systems, $(\mu,\nu)$-dichotomy, admissibility, persistence
2020 Mathematics Subject Classification:
Primary: 34D09, 39A06; Secondary: 37D25
1. Introduction
Hyperbolicity is one of the cornerstone notions in the modern theory of dynamical systems. Roughly speaking, a system is said to be hyperbolic if the phase space splits into two complementary directions such that along one of these directions we have exponential expansion with time, while in the other one we have exponential contraction. On the other hand, sometimes it may be very complicated to verify if a given system is hyperbolic. Consequently, an important problem consists in presenting different characterizations of this property.
In the present paper we are interested in characterizing the nonautonomous version of hyperbolicity. More precisely, we will present a characterization of the notion of $(\mu,\nu)$-dichotomy. The notion of $(\mu,\nu)$-dichotomy requires that the phase space splits (at each moment of time) into two directions: the stable and the unstable direction. Along stable/unstable direction dynamics contracts/expands with the rate of contraction/expansion given by a function $\mu$, while the speed of contraction/expansion is measured using a function $\nu$. We stress that the notion of $(\mu,\nu)$-dichotomy includes the notion of (nonuniform) exponential dichotomy as a very particular case. The type of characterization we are looking for has a long history, which goes back to the work of Perron [36] for ODEs, and is given in terms of what is nowadays called the admissibility property.
Given a nonautonomous dynamical system
$x_{n+1}=A_{n}x_{n},\quad n\in I,$ (1)
where $A_{n}\colon X\to X$, $n\in I$, are linear maps acting on a Banach space $(X,\|\cdot\|)$ and $I$ is an interval of ${\mathbb{Z}}$, we say that the pair $(Y,Z)$ is (properly) admissible for Eq. (1), where $Y$ and $Z$ are subspaces of $X^{I}$, if for every sequence $(y_{n})_{n\in I}$ in $Y$ there exists a (unique) sequence $(x_{n})_{n\in I}$ in $Z$ such that
$x_{n+1}=A_{n}x_{n}+y_{n+1},\quad\text{ for all }n\in I.$
Thus, a prototype result says that for $I={\mathbb{Z}}$, if $Y=Z=\ell^{\infty}$ (the space of all bounded two-sided sequences in $X$) then the proper admissibility of the pair $(Y,Z)$ is equivalent to the fact that Eq. (1) admits an exponential dichotomy (see [22]) and the characterizations of $(\mu,\nu)$-dichotomy that we are seeking have this flavor. Nevertheless, our results involve two or three (depending on $I$) admissible pairs of some special weighted spaces instead of just $\ell^{\infty}$ (see Sections 3 and 4).
The importance of our results stems from our general framework. More precisely, we are able to treat in a unified manner various settings in which no similar result has been previously obtained as well as to recover and refine several known results. We emphasize that our results hold without any bounded growth assumption for (1) and the statements make no use of Lyapunov norms. We explain a bit more the importance of these facts bellow. Moreover, as a consequence of our characterization, we obtain that the notion of $(\mu,\nu)$-dichotomy persists under small but very general linear perturbations.
1.1. Relations with previous results
As already emphasized, characterizations of dichotomies in terms of admissibility properties have a long history. For instance, a fundamental contribution to this line of research is due to Massera and Schäffer [28, 29] (see also Coppel [12]) who presented a complete characterization (in terms of admissibility) of the notion of an exponential dichotomy, extending the original work of Perron [36] (see also [24] for related results in the case of discrete time) which dealt with exponentially stable systems. The case of infinite-dimensional dynamics were first considered by Daleckiĭ and Kreĭn [13] in the case of continuous time and by Coffman and Schäffer [11] as well as Henry [22] in the case of noninvertible discrete time dynamics. More recent contributions devoted to the characterization of uniform exponential dichotomies include (but are not limited to) [1, 23, 25, 31, 32, 39, 42, 43]. We also refer to [5] for a detailed overview of this line of research and additional references.
In the case of nonuniform exponential dichotomies, there are essentially three type of results available. Firstly, there are results which do not give a complete characterization of nonuniform exponential dichotomies via admissibility but rather only sufficient conditions for the existence of dichotomy (see [30, 38, 40, 41] and references therein). Secondly, there are various results in which a complete characterization of nonuniform dichotomies via admissibility is obtained in which the corresponding input-output spaces are constructed in terms of Lyapunov norms which are used to transform nonuniform behavior into the uniform one (see [2, 3, 4, 26, 27, 48]). Finally, since Lyapunov norms are difficult to construct without firstly establishing the existence of nonuniform behavior, it was of interest to explore whether nonuniform exponential dichotomies can be characterized in terms of admissibility without the use of Lyapunov norms. The number of very recent results show that this is indeed possible (see [20, 21, 47, 49]).
As already mentioned, besides exponential dichotomies, it of interest to study dichotomic behavior when the rates of expansion/contraction along the unstable/stable direction are not of exponential type. To the best of our knowledge such dichotomies were first studied by Muldowney [34] and Naulin and Pinto [35]. More recently, a systematic study of such behavior was initiated by Barreira and Valls [6], as well as Bento and Silva [7, 8]. Dragičević [14] characterized nonuniform polynomial dichotomies in terms of admissibility for discrete dynamics (see [15] for related result for continuous time). An alternative approach which relies on the relationship between exponential and polynomial dichotomies was proposed in [17]. In the case of continuous time, a very general class of dichotomies associated to differentiable growth rates was characterized via admissibility in [16]. Finally, the most general results in the case of discrete time were obtained by Silva [44]. We stress that all these works rely on the usage of Lyapunov norms.
In the present paper we aim to characterize $(\mu,\nu)$-dichotomies in terms of admissibility without the use of Lyapunov norms. Therefore, our results complement those obtained by Silva [44]. We emphasize that even in the case of uniform dichotomies associated to a growth rate $\mu$, our results do not coincide with those in [44] as the input-output spaces are different. Moreover, in contrast to [44] we:
  1. (1)
    do not impose any bounded growth conditions;
  2. (2)
    are able to treat the case of arbitrary growth rates, while in [44] it is required that $\mu$ is “slowly growing”;
  3. (3)
    discuss the case of two-sided dynamics which was not treated in [44].
Our proof was inspired by the work of Dragičević, Zhang and Zhou [20, 21] in which the authors completely characterize, in terms of admissibility, the notion of nonuniform exponential dichotomy, which, as already mentioned, is a particular example of $(\mu,\nu)$-dichotomy.
The paper is organized as follows. In Section 2, we introduce the notion of $(\mu,\nu)$-dichotomy and the weighted spaces we are going to work with along the text. Section 3 is devoted to present the characterization as well as the robustness of $(\mu,\nu)$-dichotomies in the case of one-sided dynamics while Section 4 is devoted to study the two-sided case.
2. Preliminaries
[…]
$x_{n+1}=A_{n}x_{n},\quad n\in I.$ (2)
For $m,n\in I$, the evolution operator associated to (2) is given by
${\mathcal{A}}(m,n)=\begin{cases}A_{m-1}\cdots A_{n}&\text{for $m>n$;}\\ \text{Id}&\text{for $m=n$,}\\ \end{cases}$
where Id denotes the identity operator on $X$.
2.1. Growth rates and $(\mu,\nu)$-dichotomy
Let $\mu=(\mu_{n})_{n\in I}$ be a strictly increasing sequence of positive numbers such that
$\lim_{n\to+\infty}\mu_{n}=+\infty.$ (3)
Moreover, in the case when $I={\mathbb{Z}}$ we assume also that $\lim\limits_{n\to-\infty}\mu_{n}=0$. We call such sequence $\mu$ a growth rate. Furthermore, let $\nu=(\nu_{n})_{n\in I}$ be an arbitrary sequence with $\nu_{n}\geq 1$ for every $n\in I$.
Definition 2.1.
We say that (2) admits a $(\mu,\nu)$-dichotomy if the following conditions are satisfied:
  1. (1)
    there exists a family of projections $P_{n}$, $n\in I$, such that
    $A_{n}P_{n}=P_{n+1}A_{n};$ (4)
  2. (2)
    $A_{n}\rvert_{\operatorname{Ker}P_{n}}\colon\operatorname{Ker}P_{n}\to\operatorname{Ker}P_{n+1}$ is an invertible operator for each $n\in I$;
  3. (3)
    there exist $D,\lambda>0$ such that
    $\|{\mathcal{A}}(m,n)P_{n}\|\leq D\nu_{n}\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\lambda}\quad\text{ for }m\geq n$ (5)
    and
    $\|{\mathcal{A}}(m,n)(\text{Id}-P_{n})\|\leq D\nu_{n}\left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\lambda}\quad\text{ for }m\leq n$ (6)
    where
    ${\mathcal{A}}(m,n):=\big{(}{\mathcal{A}}(n,m)\rvert_{\operatorname{Ker}P_{m}}\big{)}^{-1}\colon\operatorname{Ker}P_{n}\to\operatorname{Ker}P_{m},$
    for $m\leq n$.
Remark 2.2.
We observe that the notion of $(\mu,\nu)$-dichotomy, which appeared earlier, for instance, in [7, 8, 37], generalizes several well-known notions of dichotomy. To illustrate this claim, let us restrict ourselves to the case when $I={\mathbb{N}}$ and suppose initially that $\nu_{n}=C$ for some $C\geq 1$ and $n\in{\mathbb{N}}$. Then, by taking $\mu_{n}=e^{n}$, $n\in{\mathbb{N}}$, we recover the notion of exponential dichotomy (see [12, 28, 29]); by taking $\mu_{n}=1+n$, $n\in{\mathbb{N}}$, we recover the notion of polynomial dichotomy (see [14]); by taking $\mu_{n}=\log(2+n)$, $n\in{\mathbb{N}}$, we recover the notion of logarithmic dichotomy (see [44]). In all these cases, if we take $\nu_{n}=\mu_{n}^{\varepsilon}$ instead of $\nu_{n}=C$ for some small $\varepsilon>0$ and $n\in{\mathbb{N}}$, we get nonuniform versions of those dichotomies.
2.2. Weighted spaces
In order to describe our main results, we need to introduce a number of special weighted spaces. Given a growth rate $\mu=(\mu_{n})_{n\in I}$ and $\beta\in{\mathbb{R}}$, let us consider the weighted space $\ell_{\beta}^{\infty}$ which consist of all sequences $\mathbf{x}=(x_{n})_{n\in I}\subset X$ such that
$\|\mathbf{x}\|_{\infty,\beta}:=\sup_{n\in I}(\mu_{n}^{\beta}\|x_{n}\|)<+\infty.$
Then, it is not difficult to check that $\|\cdot\|_{\infty,\beta}$ is a norm in $\ell_{\beta}^{\infty}$ and $(\ell_{\beta}^{\infty},\|\cdot\|_{\infty,\beta})$ is a Banach space. Similarly, if $\nu=(\nu_{n})_{n\in I}$ is a sequence such that $\nu_{n}\geq 1$ for each $n\in I$, we consider the weighted space $\ell^{1}_{\beta}$ which consist of all sequences $\mathbf{x}=(x_{n})_{n\in I}\subset X$ such that
$\|\mathbf{x}\|_{1,\beta}:=\sum_{n\in I}\mu_{n}^{\beta}\nu_{n}\|x_{n}\|<+\infty.$
Again it is not difficult to verify that $(\ell_{\beta}^{1},\|\cdot\|_{1,\beta})$ is a Banach space. Moreover, given a closed subspace $Z\subset X$, for $j\in\{1,\infty\}$, let $\ell_{\beta,Z}^{j}$ be the space which consists of all sequences $\mathbf{x}=(x_{n})_{n\in I}\in\ell_{\beta}^{j}$ such that $x_{0}\in Z$. Then, $\ell_{\beta,Z}^{j}$ is a closed subspace of $\ell_{\beta}^{j}$ and, in particular, $(\ell_{\beta,Z}^{j},\|\cdot\|_{j,\beta})$ is a Banach space. In the particular case when $Z=\{0\}$, $\ell^{1}_{\beta,Z}$ will be denoted by $\ell^{1}_{\beta,0}$.
In the case when $I={\mathbb{Z}}$, we will also need to consider some extra spaces described as follows. Observe initially that, since $(\mu_{n})_{n\in{\mathbb{Z}}}$ is strictly increasing and $\lim_{n\to-\infty}\mu_{n}=0$ and $\lim_{n\to+\infty}\mu_{n}=+\infty$, there exists $n_{0}\in{\mathbb{Z}}$ such that $\mu_{n}<1$ for every $n<n_{0}$ and $\mu_{n}\geq 1$ for every $n\geq n_{0}$. Then, we consider the space $\ell_{\beta,|\cdot|}^{\infty}$ which consist of all sequences $\mathbf{x}=(x_{n})_{n\in\mathbb{Z}}\subset X$ such that
$\|\mathbf{x}\|_{\infty,\beta,|\cdot|}:=\max\left\{\sup_{n<n_{0}}(\mu_{n}^{|\beta|}\|x_{n}\|),\sup_{n\geq n_{0}}(\mu_{n}^{-|\beta|}\|x_{n}\|)\right\}<+\infty.$
Similarly, let $\ell^{1}_{\beta,|\cdot|}$ be the space consisting of all sequences $\mathbf{x}=(x_{n})_{n\in\mathbb{Z}}\subset X$ such that
$\|\mathbf{x}\|_{1,\beta,|\cdot|}:=\sum_{n<n_{0}}\mu_{n}^{|\beta|}\nu_{n}\|x_{n}\|+\sum_{n\geq n_{0}}\mu_{n}^{-|\beta|}\nu_{n}\|x_{n}\|<+\infty.$
Once again one can easily check that $(\ell_{\beta,|\cdot|}^{\infty},\|\cdot\|_{\infty,\beta,|\cdot|})$ and $(\ell_{\beta,|\cdot|}^{1},\|\cdot\|_{1,\beta,|\cdot|})$ are Banach spaces. Observe that, even though all the spaces defined above depend on $\mu$, $\nu$ and $I$, we do not write these dependence explicitly in order to simplify notation. The dependence will be clear from the context.
3. The case of one-sided dynamics
In this section we will restrict our attention to the case of one-sided dynamics, that is, to the case when $I={\mathbb{N}}$.
3.1. Characterization of $(\mu,\nu)$-dichotomy
We will now present a characterization of $(\mu,\nu)$-dichotomy in terms of the admissibility of the spaces $\ell^{1}_{\beta,0}$ and $\ell^{\infty}_{\beta,Z}$ for some appropriate subspace $Z\subset X$ and $\beta\in{\mathbb{R}}$.
Theorem 3.1.
Suppose that $(A_{n})_{n\in\mathbb{N}}$ admits a $(\mu,\nu)$-dichotomy with respect to projections $P_{n}$ and let $\lambda>0$ be such that (5) and (6) hold. Moreover, suppose that there exists $\varepsilon\in[0,\lambda)$ such that
$\sup_{n\in\mathbb{N}}(\mu_{n}^{-\varepsilon}\nu_{n})<+\infty.$ (7)
[…]
$x_{n+1}-A_{n}x_{n}=y_{n+1},\quad n\in\mathbb{N}.$ (8)
Proof.
[…]
$x_{n}:=\sum_{k=0}^{n}\mathcal{A}(n,k)P_{k}y_{k}-\sum_{k=n+1}^{\infty}\mathcal{A}(n,k)(\text{Id}-P_{k})y_{k},\quad n\in\mathbb{N}.$ (9)
Then,
$\begin{split}\mu_{n}^{\beta}\|x_{n}\|&\leq D\mu_{n}^{\beta}\sum_{k=0}^{n}\left(\frac{\mu_{n}}{\mu_{k}}\right)^{-\lambda}\nu_{k}\|y_{k}\|+D\mu_{n}^{\beta}\sum_{k=n+1}^{\infty}\left(\frac{\mu_{k}}{\mu_{n}}\right)^{-\lambda}\nu_{k}\|y_{k}\|\\ &\leq D\sum_{k=0}^{n}\left(\frac{\mu_{n}}{\mu_{k}}\right)^{-(\lambda-\beta)}\mu_{k}^{\beta}\nu_{k}\|y_{k}\|+D\sum_{k=n+1}^{\infty}\left(\frac{\mu_{k}}{\mu_{n}}\right)^{-(\lambda+\beta)}\mu_{k}^{\beta}\nu_{k}\|y_{k}\|,\\ \end{split}$
which implies (since $\lambda-\beta>0$ and $\lambda+\beta>0$) that
$\mu_{n}^{\beta}\|x_{n}\|\leq D\sum_{k=0}^{\infty}\mu_{k}^{\beta}\nu_{k}\|y_{k}\|,\quad n\in\mathbb{N}.$
Note also that $x_{0}\in Z$ and, consequently, $\mathbf{x}=(x_{n})_{n\in\mathbb{N}}\in\ell^{\infty}_{\beta,Z}$. Moreover, it is easy to verify that $\mathbf{x}$ satisfies  (8). Indeed, given $n\in{\mathbb{N}}$,
$\begin{split}A_{n}x_{n}&=A_{n}\sum_{k=0}^{n}{\mathcal{A}}(n,k)P_{k}y_{k}-A_{n}\sum_{k=n+1}^{\infty}{\mathcal{A}}(n,k)(\text{Id}-P_{k})y_{k}\\ &=\sum_{k=0}^{n}{\mathcal{A}}(n+1,k)P_{k}y_{k}-\sum_{k=n+1}^{\infty}{\mathcal{A}}(n+1,k)(\text{Id}-P_{k})y_{k}\\ &=\sum_{k=0}^{n+1}{\mathcal{A}}(n+1,k)P_{k}y_{k}-P_{n+1}y_{n+1}\\ &\phantom{=}-\sum_{k=n+2}^{\infty}{\mathcal{A}}(n+1,k)(\text{Id}-P_{k})y_{k}-(\text{Id}-P_{n+1})y_{n+1}\\ &=x_{n+1}-y_{n+1},\end{split}$
as claimed. It remains to establish the uniqueness of such sequence $\mathbf{x}$. Suppose that $\tilde{\mathbf{x}}=(\tilde{x}_{n})_{n\in\mathbb{N}}$ is another sequence in $\ell^{\infty}_{\beta,Z}$ that satisfies (8). Then,
$x_{n}-\tilde{x}_{n}=\mathcal{A}(n,0)(x_{0}-\tilde{x}_{0}),\quad n\in\mathbb{N}.$
Consequently, since $x_{0}-\tilde{x}_{0}\in Z=\operatorname{Ker}P_{0}$, condition (6) gives us that
$\begin{split}\mu_{0}^{-\lambda}\|x_{0}-\tilde{x}_{0}\|&\leq D\mu_{n}^{-\lambda}\nu_{n}\|x_{n}-\tilde{x}_{n}\|\\ &\leq D\mu_{n}^{-(\lambda+\beta)}\nu_{n}\|\mathbf{x}-\tilde{\mathbf{x}}\|_{\infty,\beta}\\ &\leq D\mu_{n}^{-(\lambda+\beta-\varepsilon)}\|\mathbf{x}-\tilde{\mathbf{x}}\|_{\infty,\beta}\cdot\sup_{n\in\mathbb{N}}(\mu_{n}^{-\varepsilon}\nu_{n}),\end{split}$
for $n\in\mathbb{N}$. Letting $n\to\infty$, using (7) and $\beta>-(\lambda-\varepsilon)$ yields that $x_{0}=\tilde{x}_{0}$ and thus $\mathbf{x}=\tilde{\mathbf{x}}$. The proof of the theorem is completed. ∎
We now establish the converse result.
Theorem 3.2.
Suppose that there exist a closed subspace $Z\subset X$ and $\beta>0$ such that:
  1. (1)
[…]
  2. (2)
[…]
Then, $(A_{n})_{n\in\mathbb{N}}$ admits a $(\mu,\nu)$-dichotomy.
Proof.
We will split the proof into a series of auxiliary results and, in all of them, we assume that the hypotheses of Theorem 3.2 are satisfied even though we do not write it explicitly each time.
Lemma 3.3.
[…]
Proof of Lemma 3.3.
The result follows directly from the ‘uniqueness’ in our assumptions and the simple observation that $\ell^{\infty}_{\beta,Z}\subset\ell^{\infty}_{-\beta,Z}$. ∎
Let us consider
$S(n):=\left\{v\in X:\sup_{m\geq n}\|\mathcal{A}(m,n)v\|<+\infty\right\}$ (10)
and
$U(n)=\mathcal{A}(n,0)Z.$
It is easy to check that
$A_{n}S(n)\subset S(n+1)\text{ and }A_{n}U(n)\subset U(n+1)$ (11)
for every $n\in{\mathbb{N}}$.
Lemma 3.4.
We have that
$X=S(n)\oplus U(n),\quad n\in\mathbb{N}.$ (12)
Proof of Lemma 3.4.
Take initially $n\geq 1$. Given $v\in X$, we define $\mathbf{y}=(y_{m})_{m\in\mathbb{N}}$ by $y_{n}=v$ and $y_{m}=0$ for $m\neq n$. Then, $\mathbf{y}\in\ell^{1}_{\beta,0}$. By our assumption, there exists $\mathbf{x}=(x_{m})_{m\in\mathbb{N}}\in\ell^{\infty}_{\beta,Z}$ such that
$x_{m+1}-A_{m}x_{m}=y_{m+1},\quad m\in\mathbb{N}.$
In particular,
$x_{n}-A_{n-1}x_{n-1}=v.$
Then, since $x_{0}\in Z$, we have that $A_{n-1}x_{n-1}=\mathcal{A}(n,0)x_{0}\in U(n)$. Moreover, since $x_{m}=\mathcal{A}(m,n)x_{n}$ for $m\geq n$ and $\mathbf{x}\in\ell^{\infty}_{\beta,Z}$, we conclude that $x_{n}\in S(n)$. Thus, $v\in S(n)+U(n)$.
Suppose now that $v\in S(n)\cap U(n)$. Then, there exists $w\in Z$ such that $v=\mathcal{A}(n,0)w$. We define
$x_{m}:=\mathcal{A}(m,0)w,\quad m\in\mathbb{N}.$
Then, $\mathbf{x}=(x_{m})_{m\in\mathbb{N}}\in\ell^{\infty}_{-\beta,Z}$ and (8) holds with $\mathbf{y}=0$. Therefore, using the uniqueness in our assumption, we conclude that $\mathbf{x}=0$ and $v=0$ proving that (12) holds for every $n\geq 1$.
Let us now consider the case when $n=0$. Given $v\in X$, we define sequences $\mathbf{x}=(x_{m})_{m\in\mathbb{N}}$ and $\mathbf{y}=(y_{m})_{m\in\mathbb{N}}$ given by $x_{0}=v$ and $x_{m}=0$ for $m\neq n$ and $y_{1}=-A_{0}v$ and $y_{m}=0$ for $m\neq 1$. Then,
$x_{m+1}-A_{m}x_{m}=y_{m+1},\quad m\in{\mathbb{N}}.$
Moreover, since $\mathbf{y}\in\ell^{1}_{\beta,0}$, there exists $\tilde{\mathbf{x}}=(\tilde{x}_{m})_{m\in{\mathbb{N}}}\in\ell^{\infty}_{\beta,Z}$ such that (8) holds. Then,
$x_{m}-\tilde{x}_{m}=\mathcal{A}(m,0)(v-\tilde{x}_{0}),\quad\text{ for all }m\in{\mathbb{N}}.$
Consequently, since $\mathbf{x}-\tilde{\mathbf{x}}\in\ell^{\infty}_{\beta}$, it follows that $v-\tilde{x}_{0}\in S(0)$. Therefore, since $\tilde{x}_{0}\in Z$, we conclude that $v\in S(0)+U(0)$. Finally, consider $v\in S(0)\cap U(0)$ and let $x_{m}=\mathcal{A}(m,0)v$ for $m\in{\mathbb{N}}$. Then, as in the case when $n\geq 1$, $\mathbf{x}=(x_{m})_{m\in\mathbb{N}}\in\ell^{\infty}_{-\beta,Z}$ and (8) holds with $\mathbf{y}=0$. Therefore, using the uniqueness in our assumption, we conclude that $\mathbf{x}=0$ and $v=0$. This completes the proof of the lemma. ∎
Lemma 3.5.
For every $n\in{\mathbb{N}}$, the operator $A_{n}|_{U(n)}\colon U(n)\to U(n+1)$ is an isomorphism.
Proof of Lemma 3.5.
Fix $n\in{\mathbb{N}}$. Let us start proving that $A_{n}|_{U(n)}$ is injective. Suppose there exists $v\in U(n)$ such that $A_{n}v=0$. By the definition of $U(n)$ there exists $z_{0}\in Z$ such that $v=\mathcal{A}(n,0)z_{0}$. Then, the sequence $(x_{m})_{m\in{\mathbb{N}}}$ given by
$x_{m}=\mathcal{A}(m,0)z_{0},\quad m\in{\mathbb{N}},$
belongs to $\ell^{\infty}_{\beta,Z}$ and satisfies (8) with $\mathbf{y}=0$. Consequently, by the uniqueness in our hypothesis, it follows that $x_{m}=0$ for every $m\in{\mathbb{N}}$. In particular, $v=0$ and $A_{n}|_{U(n)}$ is injective.
Now, given $x\in U(n+1)$, let $z_{0}\in Z$ be such that $x=\mathcal{A}(n+1,0)z_{0}$. Set $x^{\prime}:=\mathcal{A}(n,0)z_{0}$. Then, it follows directly from the definition that $x^{\prime}\in U(n)$ and $A_{n}x^{\prime}=x$ which proves that $A_{n}|_{U(n)}\colon U(n)\to U(n+1)$ is surjective. Consequently, $A_{n}|_{U(n)}\colon U(n)\to U(n+1)$ is an isomorphism as claimed. ∎
Lemma 3.6.
For each $n\in{\mathbb{N}}$, let $P_{n}\colon X\to S(n)$ be the projection associated with (12). There exists $D>0$ such that
$\sup_{n\in\mathbb{N}}\|P_{n}\|\leq D\nu_{n},\quad n\in\mathbb{N}.$ (13)
Proof of Lemma 3.6.
[…]
$x_{n+1}^{k}-A_{n}x^{k}_{n}=y^{k}_{n+1},\text{ for every }n,k\in{\mathbb{N}}.$
Thus, letting $k\to+\infty$ we conclude that
$x_{n+1}-A_{n}x_{n}=y_{n+1},\text{ for every }n\in{\mathbb{N}}$
which implies that $T_{\beta}(\mathbf{y})=\mathbf{x}$. In particular, $T_{\beta}$ is a closed operator and thus, according to the Closed Graph Theorem (see, e.g., [46, Theorem 4.2-I, p. 181]), $T_{\beta}$ is bounded.
Now, given $n\geq 1$ and $v\in X$, consider $\mathbf{y}=(y_{m})_{m\in\mathbb{N}}$ given by $y_{n}=v$ and $y_{m}=0$ for $m\neq n$. Then, $\mathbf{y}\in\ell^{1}_{\beta,0}$. Considering $T_{\beta}(\mathbf{y})=\mathbf{x}=(x_{n})_{n\in{\mathbb{N}}}$, it follows by the proof of Lemma 3.4 that $P_{n}v=x_{n}$. Thus,
$\mu_{n}^{\beta}\|P_{n}v\|=\mu_{n}^{\beta}\|x_{n}\|\leq\|\mathbf{x}\|_{\infty,\beta}=\|T_{\beta}(\mathbf{y})\|_{\infty,\beta}\leq\|T_{\beta}\|\cdot\|\mathbf{y}\|_{1,\beta}=\|T_{\beta}\|\mu_{n}^{\beta}\nu_{n}\|v\|.$
This yields that (13) holds with $D=\|T_{\beta}\|$ for $n\geq 1$. On the other hand, (13) holds with $D=\|P_{0}\|$ for $n=0$ (recall that $\nu_{0}\geq 1$). We conclude that (13) holds with
$D=\max\left\{\|T_{\beta}\|,\|P_{0}\|\right\}>0.$
Lemma 3.7.
There exists $C>0$ such that
$\|\mathcal{A}(m,n)v\|\leq C\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}\|v\|,$
for every $m\geq n$ and $v\in S(n)$.
Proof of Lemma 3.7.
Let $n\geq 1$ and $v\in S(n)$. We define sequences $(y_{m})_{m\in\mathbb{N}}$ and $(x_{m})_{m\in\mathbb{N}}$ by
$y_{m}=\begin{cases}v&m=n;\\ 0&m\neq n,\end{cases}\quad\text{and}\quad x_{m}=\begin{cases}\mathcal{A}(m,n)v&m\geq n;\\ 0&m<n.\end{cases}$
[…]
$\mu_{m}^{\beta}\|x_{m}\|\leq\|\mathbf{x}\|_{\infty,\beta}=\|T_{\beta}(\mathbf{y})\|_{\infty,\beta}\leq K\|y\|_{1,\beta}=K\mu_{n}^{\beta}\nu_{n}\|v\|$
for $m\geq n$. Thus,
$\|\mathcal{A}(m,n)v\|\leq K\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}\|v\|$
for every $m\geq n$ and $n\geq 1$.
Now, for the case when $n=0$, if $m=n=0$ then the same estimate as before holds since $K\geq 1$. Take $m\geq 1$. Then, given $v\in S(0)$, by (11) we have that $A_{0}v\in S(1)$. Thus, using the estimate obtained above we get that
$\begin{split}\|\mathcal{A}(m,0)v\|&=\|\mathcal{A}(m,1)A_{0}v\|\leq K\left(\frac{\mu_{m}}{\mu_{1}}\right)^{-\beta}\nu_{1}\|A_{0}v\|\\ &\leq K\|A_{0}\|\frac{\nu_{1}}{\nu_{0}}\left(\frac{\mu_{0}}{\mu_{1}}\right)^{-\beta}\left(\frac{\mu_{m}}{\mu_{0}}\right)^{-\beta}\nu_{0}\|v\|.\end{split}$
Therefore, taking $C=\max\{K,K\nu_{1}\nu_{0}^{-1}\|A_{0}\|\mu_{0}^{-\beta}\mu_{1}^{\beta}\}$ we get the desired result. ∎
Lemma 3.8.
There exists $\tilde{C}>0$ such that
$\|\mathcal{A}(m,n)v\|\leq\tilde{C}\left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}\|v\|,$
for every $m\leq n$ and $v\in U(n)$.
Proof of Lemma 3.8.
Take $n\geq 1$ and $v\in U(n)$. Then by Lemma 3.5 we may consider
$x_{m}=\begin{cases}\mathcal{A}(m,n)v&m<n;\\ 0&m\geq n,\end{cases}\quad\text{and}\quad y_{m}=\begin{cases}-v&m=n;\\ 0&m\neq n.\end{cases}$
[…]
$\mu_{m}^{-\beta}\|x_{m}\|\leq\|\mathbf{x}\|_{\infty,-\beta}=\|T_{-\beta}(\mathbf{y})\|_{\infty,-\beta}\leq\tilde{C}\|y\|_{1,-\beta}=\tilde{C}\mu_{n}^{-\beta}\nu_{n}\|v\|,$
for every $m<n$. This easily implies that
$\|\mathcal{A}(m,n)v\|\leq\tilde{C}\left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}\|v\|,$
for every $m<n$ as claimed. Finally, the cases when $m=n$ or $n=0$ trivially holds since $\tilde{C}\geq 1$. ∎
For $m,n\in\mathbb{N}$, set
$\mathcal{G}(m,n):=\begin{cases}\mathcal{A}(m,n)P_{n}&m\geq n;\\ -\mathcal{A}(m,n)(\text{Id}-P_{n})&m<n.\end{cases}$
Then, by Lemmas 3.6, 3.7 and 3.8 there exists $C^{\prime}>0$ such that
$\|\mathcal{G}(m,n)\|\leq C^{\prime}\begin{cases}\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}^{2}&m\geq n;\\ \left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}^{2}&m<n.\end{cases}$ (14)
The purpose of the following lemma is to replace $\nu_{n}^{2}$ by $\nu_{n}$ in (14).
Lemma 3.9.
There exists $C^{\prime\prime}>0$ such that
$\|\mathcal{G}(m,n)\|\leq C^{\prime\prime}\begin{cases}\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}&m\geq n;\\ \left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}&m<n.\end{cases}$
Proof of the Lemma 3.9.
Take $n\geq 1$ and $v\in X$. We define a sequence $\mathbf{y}=(y_{m})_{m\in\mathbb{N}}$ by $y_{n}=v$ and $y_{m}=0$ for $m\neq n$. Then, $\mathbf{y}\in\ell_{\beta,0}^{1}\cap\ell_{-\beta,0}^{1}$. Set
$x_{m}:=\mathcal{G}(m,n)v,\quad m\in\mathbb{N}.$
It is easy to verify that (8) holds. We claim that $\mathbf{x}=(x_{m})_{m\in\mathbb{N}}\in\ell_{\beta,Z}^{\infty}$. Indeed, for $m\geq n$ we have that
$\|x_{m}\|=\|\mathcal{G}(m,n)v\|\leq C^{\prime}\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}^{2}\|v\|,$
which implies that
$\sup_{m\in\mathbb{N}}(\mu_{m}^{\beta}\|x_{m}\|)<+\infty.$
Consequently, $\mathbf{x}=(x_{m})_{m\in\mathbb{N}}\in\ell_{\beta,Z}^{\infty}$ and $T_{\beta}(\mathbf{y})=\mathbf{x}$, where $T_{\beta}$ is as in the proof of Lemma 3.6. Therefore, for $m\geq n$ we have that
$\mu_{m}^{\beta}\|\mathcal{G}(m,n)v\|=\mu_{m}^{\beta}\|x_{m}\|\leq\|\mathbf{x}\|_{\infty,\beta}\leq\|T_{\beta}\|\cdot\|\mathbf{y}\|_{1,\beta}=\|T_{\beta}\|\mu_{n}^{\beta}\nu_{n}\|v\|,$
and thus
$\|\mathcal{G}(m,n)\|\leq\|T_{\beta}\|\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n},\quad m\geq n\geq 1.$
We now consider the case when $n=0$. For $m>0$ we have
$\begin{split}\|\mathcal{G}(m,0)\|=\|\mathcal{G}(m,1)A_{0}\|&\leq\|T_{\beta}\|\cdot\|A_{0}\|\left(\frac{\mu_{m}}{\mu_{1}}\right)^{-\beta}\nu_{1}\\ &=\|T_{\beta}\|\cdot\|A_{0}\|\frac{\nu_{1}}{\nu_{0}}\left(\frac{\mu_{0}}{\mu_{1}}\right)^{-\beta}\left(\frac{\mu_{m}}{\mu_{0}}\right)^{-\beta}\nu_{0}.\end{split}$
Hence, taking
$C^{\prime\prime}\geq\max\left\{\|T_{\beta}\|,\|T_{\beta}\|\cdot\|A_{0}\|\frac{\nu_{1}}{\nu_{0}}\left(\frac{\mu_{0}}{\mu_{1}}\right)^{-\beta},\|P_{0}\|\right\}$
we get that
$\|\mathcal{G}(m,n)\|\leq C^{\prime\prime}\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n},$
for $m\geq n$. Similarly, one can treat the case when $m<n$. ∎
Now the proof of Theorem 3.2 can be easily completed by combining the previous auxiliary results. ∎
3.2. An alternative characterization of $(\mu,\nu)$-dichotomy for compact operators
We now present an alternative to Theorem 3.2 in the case when at least one of the operators $A_{n}$ is compact. For this purpose, given $\beta\in\mathbb{R}$, let
$S_{\beta}(0)=\left\{v\in X:\ \sup_{n\in\mathbb{N}}(\mu_{n}^{\beta}\|{\mathcal{A}}(n,0)v\|)<+\infty\right\}.$
Note that $S_{\beta}(0)$ is a subspace of $X$.
Theorem 3.10.
Suppose that $A_{n}$ is compact for some $n\in\mathbb{N}$ and that there exists $\beta>0$ such that:
  1. (1)
    $S_{0}(0)=S_{\beta}(0)$;
  2. (2)
[…]
  3. (3)
[…]
Then, $(A_{n})_{n\in\mathbb{N}}$ admits a $(\mu,\nu)$-dichotomy.
Proof.
For $n\in\mathbb{N}$, let $S(n)$ be given by (10). We first claim that
$A_{n}^{-1}(S(n+1))=S(n),\quad n\in\mathbb{N}.$
To this end, take $v\in A_{n}^{-1}(S(n+1))$. Then, $A_{n}v\in S(n+1)$ and thus
$\sup_{m\geq n+1}\|{\mathcal{A}}(m,n+1)A_{n}v\|=\sup_{m\geq n+1}\|{\mathcal{A}}(m,n)v\|<+\infty,$
which implies that $\sup_{m\geq n}\|{\mathcal{A}}(m,n)v\|<+\infty$, i.e. $v\in S(n)$. The converse inclusion can be established in a similar manner.
Next, we claim that for $m\geq n$,
$X={\mathcal{A}}(m,n)X+S(m).$
One can easily see that it is sufficient to consider the case when $n=0$. Then, if $m=0$, there is nothing to show since ${\mathcal{A}}(m,n)X=X$. Let us now consider the case when $m>0$. Take $v\in X$ and define $\mathbf{y}=(y_{k})_{k\in\mathbb{N}}$ by $y_{m}=v$ and $y_{k}=0$ for $k\neq m$. Then, $\mathbf{y}\in\ell_{\beta,0}^{1}$. Consequently, there exists $\mathbf{x}=(x_{k})_{k\in\mathbb{N}}\in\ell_{\beta}^{\infty}$ such that (8) holds. In particular,
$x_{m}-A_{m-1}x_{m-1}=v$
and
$x_{k}=A_{k-1}x_{k-1},\quad\text{for }k\neq m.$
Hence, since $\mathbf{x}\in\ell_{\beta}^{\infty}$, we get that $x_{m}\in S(m)$. On the other hand, $A_{m-1}x_{m-1}={\mathcal{A}}(m,0)x_{0}$, yielding that
$v=x_{m}-A_{m-1}x_{m-1}\in S(m)+{\mathcal{A}}(m,0)X.$
Thus, the desired claim holds.
We proceed by noting that each $S(m)$ is an image of a Banach space under the action of a bounded linear operator. Indeed, let $\mathcal{C}(m)$ denote the space of all sequences $\mathbf{x}=(x_{n})_{n\geq m}\subset X$ such that
$\|\mathbf{x}\|_{\mathcal{C}(m)}:=\sup_{n\geq m}\|x_{n}\|<+\infty.$
Then, $(\mathcal{C}(m),\|\cdot\|_{\mathcal{C}(m)})$ is a Banach space. Set $\mathcal{C}^{\prime}(m)$ to be the set of all sequences $\mathbf{x}=(x_{n})_{n\geq m}\in\mathcal{C}(m)$ satisfying
$x_{n+1}=A_{n}x_{n},\quad n\geq m.$
It is straightforward to verify that $\mathcal{C}^{\prime}(m)$ is a closed subspace of $\mathcal{C}(m)$ and therefore also a Banach space. Then, we observe that $S(m)=\Phi(\mathcal{C}^{\prime}(m))$, where $\Phi\colon\mathcal{C}(m)\to X$ is a bounded linear operator given by
$\Phi(\mathbf{x})=x_{m},\quad\mathbf{x}=(x_{n})_{n\geq m}\in\mathcal{C}(m).$
It follows now from [45, Lemma 3.4.] that $S(0)=S_{0}(0)$ is closed and complemented in $X$. Therefore, there exists a closed subspace $Z\subset X$ such that
$X=S(0)\oplus Z.$
Take an arbitrary $\mathbf{y}=(y_{n})_{n\in\mathbb{Z}}\in\ell_{\beta,0}^{\infty}$. By our assumption, there exists $\mathbf{x}=(x_{n})_{n\in\mathbb{N}}\in\ell_{\beta}^{\infty}$ such that (8) holds. Write $x_{0}=v_{1}+v_{2}$ with $v_{1}\in S(0)$ and $v_{2}\in Z$. Set
$\tilde{x}_{n}:=x_{n}-{\mathcal{A}}(n,0)v_{1}.$
Since $v_{1}\in S(0)=S_{0}(0)=S_{\beta}(0)$, we easily get that $\tilde{\mathbf{x}}=(\tilde{x}_{n})_{n\in\mathbb{N}}\in\ell_{\beta}^{\infty}$. Moreover, $\tilde{x}_{0}=v_{2}\in Z$. Therefore, $\tilde{\mathbf{x}}\in\ell_{\beta,Z}^{\infty}$ and it is straightforward to verify that the pair $(\tilde{\mathbf{x}},\mathbf{y})$ satisfies (8). In addition, let us suppose that the pair $(\overline{\mathbf{x}},\mathbf{y})$ satisfies (8) with $\overline{\mathbf{x}}=(\overline{x}_{n})_{n\in\mathbb{N}}\in\ell_{\beta,Z}^{\infty}$. Then, we have that $\overline{x}_{0}-\tilde{x}_{0}\in S(0)\cap Z$, which gives that $\overline{x}_{0}-\tilde{x}_{0}=0$. Therefore, $\overline{\mathbf{x}}=\tilde{\mathbf{x}}$. We have thus proved that the first assumption in the statement of Theorem 3.2 holds. Similarly, one can establish the second assumption in Theorem 3.2 (for this it is sufficient to observe that for $v\in S(0)$, $n\mapsto{\mathcal{A}}(n,0)v\in\ell_{-\beta}^{\infty}$). The desired conclusion now follows from Theorem 3.2. ∎
Remark 3.11.
  • It is easy to verify that under the assumptions of Theorem 3.1 we have that $S_{0}(0)=S_{\beta}(0)$, making this a reasonable assumption in the statement of Theorem 3.10;
  • in contrast to Theorem 3.2, the admissibility assumptions in Theorem 3.10 do not require uniqueness.
3.3. Persistence of $(\mu,\nu)$-dichotomy
In this section, as a consequence of the characterization of $(\mu,\nu)$-dichotomy given in Section 3.1, we are going to show that the notion of $(\mu,\nu)$-dichotomy persists under small linear perturbations.
Theorem 3.12.
Let $(\gamma_{n})_{n\in{\mathbb{N}}}$ be a sequence of numbers with $\gamma_{n}>0$ for every $n\in{\mathbb{N}}$ such that
$\sum_{n=0}^{\infty}\gamma_{n}<+\infty.$ (15)
Suppose that (2) admits a $(\mu,\nu)$-dichotomy with $\lambda>0$ and condition (7) is satisfied and take $\beta\in(0,\lambda-\varepsilon)$. Moreover, let $(B_{n})_{n\in{\mathbb{N}}}$ be a sequence of operators in $\mathcal{B}(X)$ with the property that there exists $c>0$ such that
$\|B_{n}\|\leq\frac{c\gamma_{n}\mu_{n}^{\beta}}{\nu_{n+1}\mu_{n+1}^{\beta}},\quad\text{for $n\in\mathbb{N}$.}$ (16)
Then, if $c$ is sufficiently small we have that the nonautonomous difference equation
$x_{n+1}=(A_{n}+B_{n})x_{n},\quad n\in{\mathbb{N}}$ (17)
also admits a $(\mu,\nu)$-dichotomy.
Remark 3.13.
Observe that in the case when $\mu_{n+1}\leq K\mu_{n}$ for every $n\in{\mathbb{N}}$ and some $K>0$, that is, the sequence $(\mu_{n})_{n\in{\mathbb{N}}}$ does not grow faster than exponential, condition (16) can be reduced to
$\|B_{n}\|\leq\frac{c\gamma_{n}}{\nu_{n+1}},$ (18)
for $n\in{\mathbb{N}}$ and $c>0$ small enough. In particular, this comment applies to the classical settings of exponential, polynomial and logarithmic dichotomy. Observe moreover that whenever we are in the case of an uniform dichotomy, meaning that the sequence $(\nu_{n})_{n\in{\mathbb{N}}}$ is constant, condition (18) is basically saying that the classical $\ell^{1}$-norm of $(\|B_{n}\|)_{n\in{\mathbb{N}}}$ is small. This indicates that our robustness is, in general, not optimal as, it is for example well-known, that the notion of uniform exponentially dichotomy persists under the requirement that $\sup_{n}\|B_{n}\|$ is small. On the other hand, in the case of polynomial and logarithmic dichotomy it does generalize existing results [14, 17, 18, 44] since, for instance, we do not impose any bounded growth condition in (2) contrary to what is done in those works.
We note that the robustness of $(\mu,\nu)$-dichotomies was discussed in [10, Theorem 3.9]. There are important differences between this result and our Theorem 3.12. Firstly, in [10] only the case of invertible dynamics is discussed. Moreover, [10, Theorem 3.9] is not applicable in the case where $\nu_{n}=1$ for each $n$, since in this setting [10, (2.5)] is not fulfilled. In particular, it is not applicable to the classical settings of uniform exponential, polynomial and logarithmic dichotomy.
Finally, we would like to compare Theorem 3.12 with an unpublished work [9]. Letting $a_{n}=\mu_{n}^{\lambda}$ and $b_{n}=\mu_{n}^{-\lambda}$, we observe that [9, Corollary 6] gives the same conclusion as Theorem 3.12 under a slightly stronger condition:
$\|B_{n}\|\leq\frac{c\gamma_{n}\mu_{n}^{\lambda}}{\nu_{n+1}\mu_{n+1}^{\lambda}},\quad\text{for $n\in\mathbb{N}$,}$ (19)
where $c>0$ is sufficiently small and $(\gamma_{n})$ satisfying (15). Indeed, observe that since $\beta<\lambda$ and $\mu_{n}/\mu_{n+1}<1$ one has
$\left(\frac{\mu_{n}}{\mu_{n+1}}\right)^{\lambda}\leq\left(\frac{\mu_{n}}{\mu_{n+1}}\right)^{\beta}.$
Consequently, (19) implies (16), while the converse is not true (if $(\mu_{n})_{n}$ is not slowly varying). Therefore, our result is slightly stronger than [9, Corollary 6]. We stress that the arguments in [9] and the present paper are completely different. In particular, the results in [9] do not rely on admissibility.
In order to prove Theorem 3.12 let us introduce some terminology. Given a closed subspace $Z\subset X$ and $\beta>0$, let us consider consider the linear operators $\mathbf{A}_{\beta}\colon\mathcal{D}(\mathbf{A}_{\beta})\subset\ell^{\infty}_{\beta,Z}\to\ell^{1}_{\beta,0}$ and $\mathbf{A}_{-\beta}\colon\mathcal{D}(\mathbf{A}_{-\beta})\subset\ell^{\infty}_{-\beta,Z}\to\ell^{1}_{-\beta,0}$ given by
$(\mathbf{A}_{j}\mathbf{x})_{n}=x_{n}-A_{n-1}x_{n-1}\text{ for }n\geq 1$
and $(\mathbf{A}_{j}\mathbf{x})_{0}=0$ for $j=\beta,-\beta$ where $\mathcal{D}(\mathbf{A}_{\beta})=\{\mathbf{x}\in\ell^{\infty}_{\beta,Z}:\mathbf{A}_{\beta}\mathbf{x}\in\ell^{1}_{\beta,0}\}$ and $\mathcal{D}(\mathbf{A}_{-\beta})=\{\mathbf{x}\in\ell^{\infty}_{-\beta,Z}:\mathbf{A}_{-\beta}\mathbf{x}\in\ell^{1}_{-\beta,0}\}$. Then, using this terminology we can reformulate Theorems 3.1 and 3.2 as follows.
Theorem 3.14.
Let us consider the following conditions:
  • i)
    Eq. (2) admits a $(\mu,\nu)$-dichotomy.
  • ii)
    There exist a closed subspace $Z\subset X$ and $\beta>0$ such that for every $\mathbf{y}^{1}\in\ell^{1}_{\beta,0}$ there exists a unique $\mathbf{x}^{1}\in\ell^{\infty}_{\beta,Z}$ satisfying $\mathbf{A}_{\beta}\mathbf{x}^{1}=\mathbf{y}^{1}$ and for every $\mathbf{y}^{2}\in\ell^{1}_{-\beta,0}$ there exists a unique $\mathbf{x}^{2}\in\ell^{\infty}_{-\beta,Z}$ satisfying $\mathbf{A}_{-\beta}\mathbf{x}^{2}=\mathbf{y}^{2}$. In other words, operators $\mathbf{A}_{\beta}$ and $\mathbf{A}_{-\beta}$ are invertible.
Thus, if $i)$ holds and (7) is satisfied then $ii)$ also holds. Reciprocally, if $ii)$ holds then $i)$ also holds.
Proof of Theorem 3.12.
Let $Z\subset X$ and $\beta>0$ be given by Theorem 3.14. From Theorem 3.1 it follows that $\beta$ may be any value in $(0,\lambda-\varepsilon)$. In particular, we may assume without loss of generality that $\beta$ is such that (16) is satisfied. Let us also consider the operators $\mathbf{A}_{\beta}$ and $\mathbf{A}_{-\beta}$ given above. We endow $\mathcal{D}(\mathbf{A}_{\beta})$ and $\mathcal{D}(\mathbf{A}_{-\beta})$ with the graph norms
$\|\mathbf{x}\|_{\mathbf{A}_{\beta}}=\|\mathbf{x}\|_{\infty,\beta}+\|\mathbf{A}_{\beta}(\mathbf{x})\|_{1,\beta}$
and
$\|\mathbf{x}\|_{\mathbf{A}_{-\beta}}=\|\mathbf{x}\|_{\infty,-\beta}+\|\mathbf{A}_{-\beta}(\mathbf{x})\|_{1,-\beta},$
respectively. By proceeding as in the proof of Lemma 3.6 we can conclude that $\mathbf{A}_{\beta}$ and $\mathbf{A}_{-\beta}$ are closed operators. In particular, $(\mathcal{D}(\mathbf{A}_{j}),\|\cdot\|_{\mathbf{A}_{j}})$, $j=\beta,-\beta$, are Banach spaces. Consider also the operators $\mathbf{B}_{\beta}\colon\ell^{\infty}_{\beta,Z}\to\ell^{\infty}_{\beta,0}$ and $\mathbf{B}_{-\beta}\colon\ell^{\infty}_{-\beta,Z}\to\ell^{\infty}_{-\beta,0}$ given by
$(\mathbf{B}_{j}\mathbf{x})_{n}=B_{n-1}x_{n-1}\text{ for }n\geq 1$
and $(\mathbf{B}_{j}\mathbf{x})_{0}=0$ for $j=\beta,-\beta$. Observe that condition (16) guarantee that both $\mathbf{B}_{\beta}$ and $\mathbf{B}_{-\beta}$ are well-defined.
Now, using (16) we get that for every $\mathbf{x}\in\mathcal{D}(\mathbf{A}_{\beta})$,
$\begin{split}\|((\mathbf{A}_{\beta}-\mathbf{B}_{\beta})\mathbf{x})_{n}\|&\leq\|(\mathbf{B}_{\beta}\mathbf{x})_{n}\|+\|(\mathbf{A}_{\beta}\mathbf{x})_{n}\|\\ &\leq\frac{c\gamma_{n-1}\mu_{n-1}^{\beta}}{\nu_{n}\mu_{n}^{\beta}}\|x_{n-1}\|+\|(\mathbf{A}_{\beta}\mathbf{x})_{n}\|\end{split}$ (20)
for every $n\geq 1$. Consequently, using (15) and that $\mathbf{x}\in\ell^{\infty}_{\beta}$ and $\mathbf{A}_{\beta}\mathbf{x}\in\ell^{1}_{\beta,0}$ we get that
$\begin{split}\|(\mathbf{A}_{\beta}-\mathbf{B}_{\beta})\mathbf{x}\|_{1,\beta}&\leq\sum_{n=0}^{\infty}c\gamma_{n}\mu_{n}^{\beta}\|x_{n}\|+\|\mathbf{A}_{\beta}\mathbf{x}\|_{1,\beta}\\ &\leq c\|\mathbf{x}\|_{\infty,\beta}\sum_{n=0}^{\infty}\gamma_{n}+\|\mathbf{A}_{\beta}\mathbf{x}\|_{1,\beta}<+\infty\end{split}$ (21)
and $(\mathbf{A}_{\beta}-\mathbf{B}_{\beta})\mathbf{x}\in\ell^{1}_{\beta,0}$. In particular, the operator $\mathbf{A}_{\beta}-\mathbf{B}_{\beta}\colon(\mathcal{D}(\mathbf{A}_{\beta}),\|\cdot\|_{\mathbf{A}_{\beta}})\to\ell^{1}_{\beta,0}$ given by $(\mathbf{A}_{\beta}-\mathbf{B}_{\beta})\mathbf{x}$ is well-defined. Moreover, by (21) we have that
$\|(\mathbf{A}_{\beta}-\mathbf{B}_{\beta})\mathbf{x}\|_{1,\beta}\leq K\|\mathbf{x}\|_{\mathbf{A}_{\beta}}$
for some $K>0$ and, consequently, $\mathbf{A}_{\beta}-\mathbf{B}_{\beta}$ is bounded. Furthermore, using part of the estimate obtained in (20) we get that
$\begin{split}\|\mathbf{A}_{\beta}\mathbf{x}-(\mathbf{A}_{\beta}-\mathbf{B}_{\beta})\mathbf{x}\|_{1,\beta}\leq c\sum_{n=0}^{\infty}\gamma_{n}\|\mathbf{x}\|_{\infty,\beta}\leq c\sum_{n=0}^{\infty}\gamma_{n}\|\mathbf{x}\|_{\mathbf{A}_{\beta}}.\end{split}$
Therefore, since $\mathbf{A}_{\beta}$ is invertible (recall Theorem 3.14), we conclude that for $c>0$ small enough $\mathbf{A}_{\beta}-\mathbf{B}_{\beta}$ is also invertible. Proceeding in a similar manner we conclude that $\mathbf{A}_{-\beta}-\mathbf{B}_{-\beta}\colon(\mathcal{D}(\mathbf{A}_{-\beta}),\|\cdot\|_{\mathbf{A}_{-\beta}})\to\ell^{1}_{-\beta,0}$ is also a well-defined bounded and invertible linear operator. Then, combining these two observations with Theorem 3.14 we get that (17) admits a $(\mu,\nu)$-dichotomy. ∎
We now present an example to which our results are applicable to but those of [44] are not. Observe that in this example condition [44, (10)] is not satisfied.
Example 3.15.
Take $X=\mathbb{R}$, and consider any growth rate $\mu=(\mu_{n})_{n\in\mathbb{N}}$ such that $\lim_{n\to\infty}\frac{\mu_{n+1}}{\mu_{n}}=\infty$. For example, we can take $\mu_{n}=e^{e^{n}}$, $n\in\mathbb{N}$. Moreover, let
$A_{n}=\left(\frac{\mu_{n+1}}{\mu_{n}}\right)^{-\frac{1}{2}},\quad n\in\mathbb{N}.$
Observe that
$\mathcal{A}(m,n)=\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\frac{1}{2}}\quad\text{for $m\geq n$.}$
Consequently, the sequence $(A_{n})_{n\in\mathbb{N}}$ admits a $(\mu,\nu)$-dichotomy where $\nu_{n}=1$ for $n\in\mathbb{N}$ with projections $P_{n}=\text{Id}$ for $n\in\mathbb{N}$. Consider the sequence $\mathbf{y}=(y_{n})_{n\in\mathbb{N}}\subset\mathbb{R}$ given by $y_{0}=0$ and $y_{n}=1$ for $n\geq 1$. Clearly, $\mathbf{y}$ is bounded. We claim that there does not exist a bounded sequence $\mathbf{x}=(x_{n})_{n\in\mathbb{N}}\subset\mathbb{R}$ with $x_{0}=0$ such that
$\phi_{n+1}^{\mu}(x_{n+1}-A_{n}x_{n})=y_{n+1},\quad n\in\mathbb{N},$
where $\phi_{n}^{\mu}:=\frac{\mu_{n}}{\mu_{n+1}-\mu_{n}}$. Indeed, assuming such a sequence does exist, one can easily verify that
$x_{n}=\sum_{k=1}^{n}\frac{1}{\phi_{k}^{\mu}}\mathcal{A}(n,k)y_{k}=\sum_{k=1}^{n}\frac{1}{\phi_{k}^{\mu}}\left(\frac{\mu_{n}}{\mu_{k}}\right)^{-\frac{1}{2}}=\mu_{n}^{-\frac{1}{2}}\sum_{k=1}^{n}\frac{1}{\phi_{k}^{\mu}}\mu_{k}^{\frac{1}{2}},\quad n\geq 1.$
On the other hand, by using the first inequality in [19, (7)] (whose proof does not rely on the slow-growth property [19, p.5 (iii)]), we have that
$\sum_{k=1}^{n}\frac{1}{\phi_{k}^{\mu}}\mu_{k}^{\frac{1}{2}}=\sum_{k=1}^{n}\mu_{k}^{-\frac{1}{2}}(\mu_{k+1}-\mu_{k})\geq 2(\mu_{n+1}^{\frac{1}{2}}-\mu_{1}^{\frac{1}{2}}),$
and consequently
$x_{n}\geq 2\left((\mu_{n+1}/\mu_{n})^{\frac{1}{2}}-(\mu_{1}/\mu_{n})^{\frac{1}{2}}\right),\quad n\geq 1.$
Since $\lim_{n\to\infty}\frac{\mu_{n+1}}{\mu_{n}}=\infty$, we obtain a contradiction. This shows that [44, Theorem 3.3] does not hold in this case (see [44, Remark 2]).
Remark 3.16.
We note that the example given in [14, Example 1] illustrates that the bounded growth condition [44, (14)] cannot be omitted as an assumption in [44, Theorem 3.4]. However, our results do not require such an assumption.
4. The case of two-sided dynamics
In this section we are going to consider the case of two-sided dynamics, that is, the case when $I={\mathbb{Z}}$.
4.1. Characterization of $(\mu,\nu)$-dichotomy
Following the ideas of Section 3.1, we will now present a characterization of $(\mu,\nu)$-dichotomy in terms of the admissibility of certain weighted spaces. In the present context, in addition to the admissibility of spaces of the form $\ell^{1}_{\beta}$ and $\ell^{\infty}_{\beta}$ for some appropriate values of $\beta\in{\mathbb{R}}$, we will also have to consider the admissibility of spaces $\ell^{1}_{\beta,|\cdot|}$ and $\ell^{\infty}_{\beta,|\cdot|}$.
Theorem 4.1.
Suppose that $(A_{n})_{n\in\mathbb{Z}}$ admits a $(\mu,\nu)$-dichotomy with respect to projections $P_{n}$ and let $\lambda>0$ be such that (5) and (6) hold. Moreover, suppose that there exists $\varepsilon\in[0,\lambda)$ such that
$\sup_{n\in\mathbb{N}}(\mu_{-n}^{\varepsilon}\nu_{-n})<+\infty\;\text{ and }\;\sup_{n\in\mathbb{N}}(\mu_{n}^{-\varepsilon}\nu_{n})<+\infty.$ (22)
Then, for each $\beta\in(-(\lambda-\varepsilon),\lambda-\varepsilon)$ the following holds:
  1. (1)
[…]
    $x_{n+1}-A_{n}x_{n}=y_{n+1}\quad n\in\mathbb{Z}.$ (23)
  2. (2)
    for every $\mathbf{y}=(y_{n})_{n\in\mathbb{Z}}\in\ell^{1}_{\beta,|\cdot|}$, there exists a unique $\mathbf{x}=(x_{n})_{n\in\mathbb{N}}\in\ell_{\beta,|\cdot|}^{\infty}$ such that (23) holds.
Proof.
[…]
$x_{n}:=\sum_{k=-\infty}^{n}\mathcal{A}(n,k)P_{k}y_{k}-\sum_{k=n+1}^{\infty}\mathcal{A}(n,k)(\text{Id}-P_{k})y_{k},\quad n\in\mathbb{Z}.$ (24)
Then, proceeding as in the proof of Theorem 3.1 we obtain that $\mathbf{x}=(x_{n})_{n\in\mathbb{Z}}\in\ell^{\infty}_{\beta}$ and that (23) is satisfied. We now establish the uniqueness of such $\mathbf{x}$. For this purpose, by linearity, all we have to do is to prove that the unique sequence $\mathbf{z}=(z_{n})_{n\in\mathbb{Z}}$ in $\ell^{\infty}_{\beta}$ that satisfies
$z_{n+1}=A_{n}z_{n},\quad n\in\mathbb{Z}$
is the null sequence. That is, $z_{n}=0$ for all $n\in{\mathbb{Z}}$. Let us consider
$z^{s}_{n}=P_{n}z_{n}\;\text{ and }\;z^{u}_{n}=(\text{Id}-P_{n})z_{n}\;\text{ for }n\in{\mathbb{Z}}.$
Then, $z_{n}=z^{s}_{n}+z^{u}_{n}$ and, by (4),
$z^{s}_{n+1}=A_{n}z^{s}_{n}\;\text{ and }\;z^{u}_{n+1}=A_{n}z^{u}_{n}\;\text{ for all }n\in{\mathbb{Z}}.$
Thus, using (5), we get that for every $m\geq n$,
$\begin{split}\mu_{m}^{\lambda}\|z^{s}_{m}\|&=\mu_{m}^{\lambda}\|\mathcal{A}(m,n)z^{s}_{n}\|\\ &=\mu_{m}^{\lambda}\|\mathcal{A}(m,n)P_{n}z_{n}\|\\ &\leq D\nu_{n}\mu_{n}^{\lambda}\|z_{n}\|\\ &\leq D\nu_{n}\mu_{n}^{\lambda-\beta}\|\mathbf{z}\|_{\infty,\beta}\\ &\leq D\mu_{n}^{\lambda-\beta-\varepsilon}\|\mathbf{z}\|_{\infty,\beta}\cdot\sup_{k\leq m}(\mu_{k}^{\varepsilon}\nu_{k}).\end{split}$
Consequently, since $\lim_{n\to-\infty}\mu_{n}=0$, letting $n\to-\infty$ and using (22) and $\beta<\lambda-\varepsilon$ it follows that $z^{s}_{m}=0$ for every $m\in{\mathbb{Z}}$. Similarly, using (6), we get that for every $m<n$,
$\begin{split}\mu_{m}^{-\lambda}\|z^{u}_{m}\|&=\mu_{m}^{-\lambda}\|\mathcal{A}(m,n)z^{u}_{n}\|\\ &=\mu_{m}^{-\lambda}\|\mathcal{A}(m,n)(\text{Id}-P_{n})z_{n}\|\\ &\leq D\nu_{n}\mu_{n}^{-\lambda}\|z_{n}\|\\ &\leq D\nu_{n}\mu_{n}^{-(\lambda+\beta)}\|\mathbf{z}\|_{\infty,\beta}\\ &\leq D\mu_{n}^{-(\lambda+\beta-\varepsilon)}\|\mathbf{z}\|_{\infty,\beta}\cdot\sup_{k\geq m}(\mu_{k}^{-\varepsilon}\nu_{k}).\end{split}$
Therefore, since $\lim_{n\to+\infty}\mu_{n}=+\infty$, letting $n\to+\infty$ and using (22) and $\beta>-(\lambda-\varepsilon)$ it follows that $z^{u}_{m}=0$ for every $m\in{\mathbb{Z}}$. Combining these observations we conclude that $z_{m}=z^{s}_{m}+z^{u}_{m}=0$ for every $m\in{\mathbb{Z}}$ completing the proof of the first claim in the theorem. Let us now prove the second one.
Given $\beta\in(-(\lambda-\varepsilon),\lambda-\varepsilon)$ and $\mathbf{y}=(y_{n})_{n\in\mathbb{Z}}\in\ell^{1}_{\beta,|\cdot|}$, let us consider $x_{n}$, $n\in{\mathbb{Z}}$, defined by (24). Then, using (5) and (6), if $n\geq n_{0}$ we get that
$\begin{split}\|x_{n}\|&\leq D\sum_{k=-\infty}^{n}\left(\frac{\mu_{n}}{\mu_{k}}\right)^{-\lambda}\nu_{k}\|y_{k}\|+D\sum_{k=n+1}^{\infty}\left(\frac{\mu_{k}}{\mu_{n}}\right)^{-\lambda}\nu_{k}\|y_{k}\|\\ &=D\sum_{k=-\infty}^{n_{0}-1}\left(\frac{\mu_{n}}{\mu_{k}}\right)^{-\lambda}\nu_{k}\|y_{k}\|+D\sum_{k=n_{0}}^{n}\left(\frac{\mu_{n}}{\mu_{k}}\right)^{-\lambda}\nu_{k}\|y_{k}\|\\ &\phantom{=}+D\sum_{k=n+1}^{\infty}\left(\frac{\mu_{k}}{\mu_{n}}\right)^{-\lambda}\nu_{k}\|y_{k}\|\\ &=D\sum_{k=-\infty}^{n_{0}-1}\left(\frac{\mu_{n}}{\mu_{k}}\right)^{-(\lambda-|\beta|)}\mu_{n}^{-|\beta|}\mu_{k}^{|\beta|}\nu_{k}\|y_{k}\|\\ &\phantom{=}+D\sum_{k=n_{0}}^{n}\left(\frac{\mu_{n}}{\mu_{k}}\right)^{-(\lambda+|\beta|)}\mu_{n}^{|\beta|}\mu_{k}^{-|\beta|}\nu_{k}\|y_{k}\|\\ &\phantom{=}+D\sum_{k=n+1}^{\infty}\left(\frac{\mu_{k}}{\mu_{n}}\right)^{-(\lambda-|\beta|)}\mu_{n}^{|\beta|}\mu_{k}^{-|\beta|}\nu_{k}\|y_{k}\|.\\ \end{split}$
Thus, since $(\mu_{k})_{k\in{\mathbb{Z}}}$ is strictly increasing, $|\beta|<\lambda$ and $\mu_{n}\geq 1$ for every $n\geq n_{0}$, it follows that
$\begin{split}\|x_{n}\|&\leq D\mu_{n}^{|\beta|}\left(\sum_{k=-\infty}^{n_{0}-1}\mu_{k}^{|\beta|}\nu_{k}\|y_{k}\|+\sum_{k=n_{0}}^{\infty}\mu_{k}^{-|\beta|}\nu_{k}\|y_{k}\|\right)\\ \end{split}$
which implies that
$\mu_{n}^{-|\beta|}\|x_{n}\|\leq D\|\textbf{y}\|_{1,\beta,|\cdot|}.$
Similarly, in the case when $n<n_{0}$ we can prove that
$\mu_{n}^{|\beta|}\|x_{n}\|\leq D\|\textbf{y}\|_{1,\beta,|\cdot|}.$
Consequently, $\textbf{x}=(x_{n})_{n\in{\mathbb{Z}}}\in\ell^{\infty}_{\beta,|\cdot|}$. Moreover, proceeding again as in the proof of Theorem 3.1 we conclude that (23) is satisfied. Finally, the uniqueness of this $\mathbf{x}\in\ell^{\infty}_{\beta,|\cdot|}$ can be obtained via an argument similar to the one we did in the proof of the first claim of the theorem. The proof of the theorem is completed. ∎
Our next theorem gives us the converse result of Theorem 4.1.
Theorem 4.2.
Suppose that there exists $\beta>0$ such that:
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
    for each $\mathbf{y}=(y_{n})_{n\in\mathbb{Z}}\in\ell^{1}_{\beta,|\cdot|}$, there exists a unique $\mathbf{x}=(x_{n})_{n\in\mathbb{Z}}\in\ell_{\beta,|\cdot|}^{\infty}$ such that (23) holds.
Then, $(A_{n})_{n\in\mathbb{Z}}$ admits a $(\mu,\nu)$-dichotomy.
Proof.
We will proceed as in the proof of Theorem 3.2 and again, as in the aforementioned proof, in all the auxiliary results we assume that the hypotheses of Theorem 4.2 are satisfied even though we do not write it explicitly each time. We start with two auxiliary observations.
Lemma 4.3.
[…]
Proof of Lemma 4.3.
Since $\ell^{1}_{\beta}\subset\ell^{1}_{\beta,|\cdot|}$ and $\ell^{1}_{-\beta}\subset\ell^{1}_{\beta,|\cdot|}$, it follows that $\mathbf{y}\in\ell^{1}_{\beta,|\cdot|}$. Similarly, since $\ell^{\infty}_{\beta}\subset\ell^{\infty}_{\beta,|\cdot|}$ and $\ell^{\infty}_{-\beta}\subset\ell^{\infty}_{\beta,|\cdot|}$, it follows that $\mathbf{x}^{1},\mathbf{x}^{2}\in\ell^{\infty}_{\beta,|\cdot|}$. Thus, by the uniqueness given in the third assumption of the statement of Theorem 4.2 it follows that $x_{n}^{1}=x_{n}^{2}$ for every $n\in\mathbb{Z}$ as claimed. ∎
Let us consider $T_{j}\colon\ell^{1}_{j}\to\ell^{\infty}_{j}$ given by $T_{j}(\mathbf{y})=\mathbf{x}$ where $\mathbf{x}$ is the unique element in $l^{\infty}_{j}$ such that (23) holds for $j=\beta,-\beta$.
Lemma 4.4.
Then $T_{\beta}$ and $T_{-\beta}$ are bounded linear operators.
Proof of Lemma 4.4.
The proof can be obtained by proceeding as in the proof of Lemma 3.6. ∎
Given $n\in{\mathbb{Z}}$, let us consider
$S(n):=\left\{v\in X:\;\sup_{m\geq n}\|{\mathcal{A}}(m,n)v\|<+\infty\right\}.$
Similarly, let $U(n)$ be the space of all $v\in X$ for which there exists a sequence $(z_{m})_{m\leq n}$ such that $z_{n}=v$, $z_{m}=A_{m-1}z_{m-1}$ for every $m\leq n$ and $\sup_{m\leq n}\|z_{m}\|<+\infty$.
It is easy to see that
$A_{n}S(n)\subset S(n+1)\quad\text{ and }\quad A_{n}U(n)\subset U(n+1)$
for every $n\in{\mathbb{Z}}$. Moreover, we have the following observations.
Lemma 4.5.
For every $n\in{\mathbb{Z}}$,
$X=S(n)\oplus U(n).$ (25)
Proof of Lemma 4.5.
Fix $n\in{\mathbb{Z}}$. Given $v\in X$, we define $\mathbf{y}=(y_{m})_{m\in\mathbb{Z}}$ by $y_{n}=v$ and $y_{m}=0$ for $m\neq n$. Then, $\mathbf{y}\in\ell^{1}_{\beta}\cap\ell^{1}_{-\beta}$. By our assumption and Lemma 4.3, there exists $\mathbf{x}=(x_{m})_{m\in\mathbb{Z}}\in\ell^{\infty}_{\beta}\cap\ell^{\infty}_{-\beta}$ such that
$x_{m+1}-A_{m}x_{m}=y_{m+1},\quad m\in\mathbb{Z}.$
In particular,
$x_{n}-v=A_{n-1}x_{n-1}$ (26)
and
$x_{m+1}=A_{m}x_{m}\quad\text{ for all }m\neq n-1.$ (27)
Thus, (27) combined with the fact that $(x_{m})_{m\in{\mathbb{Z}}}\in\ell_{\beta}^{\infty}$ implies that $x_{n}\in S(n)$. Similarly, (26) and (27) combined with the fact that $(x_{m})_{m\in{\mathbb{Z}}}\in\ell_{-\beta}^{\infty}$ implies that $x_{n}-v\in U(n)$. Therefore,
$v=x_{n}+(v-x_{n})\in S(n)+U(n)$
and $X=S(n)+U(n)$.
Suppose now that $v\in S(n)\cap U(n)$. Then there exists $(z_{m})_{m\leq n}$ such that $z_{n}=v$, $z_{m}=A_{m-1}z_{m-1}$ for every $m\leq n$ and $\sup_{m\leq n}\|z_{m}\|<+\infty$. Define
$x_{m}=\left\{\begin{array}[]{ll}z_{m},&\text{ for }m\leq n;\\ {\mathcal{A}}(m,n)v,&\text{ for }m>n.\end{array}\right.$
Thus, since $v\in S(n)$ and $\sup_{m\leq n}\|z_{m}\|<+\infty$, it follows that $\sup_{m\in{\mathbb{Z}}}\|x_{m}\|<+\infty$. In particular, $(x_{m})_{m\in{\mathbb{Z}}}\subset\ell^{\infty}_{\beta,|\cdot|}$ and
$x_{m+1}-A_{m}x_{m}=0\;\text{ for every }m\in{\mathbb{Z}}.$
Therefore, from the uniqueness given in the third hypothesis of Theorem 4.2 we get that $x_{m}=0$ for every $m\in{\mathbb{Z}}$ which implies that $v=0$ and $S(n)\cap U(n)=\{0\}$. This concludes the proof. ∎
Lemma 4.6.
For every $n\in{\mathbb{Z}}$, the operator $A_{n}|_{U(n)}\colon U(n)\to U(n+1)$ is an isomorphism.
Proof of Lemma 4.6.
Fix $n\in{\mathbb{Z}}$. Let us start by observing that $A_{n}|_{U(n)}$ is injective. Suppose there exists $v\in U(n)$ such that $A_{n}v=0$. By the definition of $U(n)$ there exists $(z_{m})_{m\leq n}$ such that $z_{n}=v$, $z_{m}=A_{m-1}z_{m-1}$ for every $m\leq n$ and $\sup_{m\leq n}\|z_{m}\|<+\infty$. Then the sequence $(x_{m})_{m\in{\mathbb{Z}}}$ given by
$x_{m}=\left\{\begin{array}[]{ll}z_{m},&\text{ for }m\leq n;\\ 0,&\text{ for }m>n,\end{array}\right.$
is in $l^{\infty}_{\beta}$ and satisfies (23) with $\mathbf{y}=(y_{m})_{m\in{\mathbb{Z}}}$ and $y_{m}=0$ for every $m\in{\mathbb{Z}}$. Consequently, by the uniqueness in the hypothesis of Theorem 4.2, it follows that $x_{m}=0$ for every $m\in{\mathbb{Z}}$. In particular, $v=0$ and $A_{n}|_{U(n)}$ is injective.
Now, given $v\in U(n+1)$, let us consider a sequence $(z_{m})_{m\leq n+1}$ such that $z_{n+1}=v$, $z_{m}=A_{m-1}z_{m-1}$ for every $m\leq n+1$ and $\sup_{m\leq n+1}\|z_{m}\|<+\infty$. Then, $z_{n}\in U(n)$ and $A_{n}z_{n}=v$ which proves that $A_{n}|_{U(n)}\colon U_{n}\to U(n+1)$ is surjective. ∎
Lemma 4.7.
For each $n\in{\mathbb{Z}}$, let $P_{n}\colon X\to S(n)$ be the projection associated with (25). Then, there exists $D>0$ such that
$\|P_{n}\|\leq D\nu_{n},\quad n\in\mathbb{Z}.$ (28)
Proof of Lemma 4.7.
The proof of this result is the same, mutatis mutandis, as the proof of Lemma 3.6. ∎
Lemma 4.8.
There exists $C>0$ such that
$\|\mathcal{A}(m,n)v\|\leq C\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}\|v\|,$
for every $m\geq n$ and $v\in S(n)$.
Proof of Lemma 4.8.
Let $n\in{\mathbb{Z}}$ and $v\in S(n)$. We define sequences $(y_{m})_{m\in\mathbb{Z}}$ and $(x_{m})_{m\in\mathbb{Z}}$ by
$y_{m}=\begin{cases}v&m=n;\\ 0&m\neq n,\end{cases}\quad\text{and}\quad x_{m}=\begin{cases}\mathcal{A}(m,n)v&m\geq n;\\ 0&m<n.\end{cases}$
[…]
$\mu_{m}^{\beta}\|x_{m}\|\leq\|\mathbf{x}\|_{\infty,\beta}=\|T_{\beta}(\mathbf{y})\|_{\infty,\beta}\leq C\|y\|_{1,\beta}=C\mu_{n}^{\beta}\nu_{n}\|v\|$
for $m\geq n$. Thus,
$\|\mathcal{A}(m,n)v\|\leq C\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}\|v\|$
for every $m\geq n$ as claimed. ∎
Lemma 4.9.
There exists $\tilde{C}>0$ such that
$\|\mathcal{A}(m,n)v\|\leq\tilde{C}\left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}\|v\|,$
for every $m\leq n$ and $v\in U(n)$.
Proof of Lemma 4.9.
Take $n\in{\mathbb{Z}}$ and $v\in U(n)$. Then by Lemma 4.6 we may consider
$y_{m}=\begin{cases}-v&m=n;\\ 0&m\neq n,\end{cases}\quad\text{and}\quad x_{m}=\begin{cases}\mathcal{A}(m,n)v&m<n;\\ 0&m\geq n.\end{cases}$
[…]
$\mu_{m}^{-\beta}\|x_{m}\|\leq\|\mathbf{x}\|_{\infty,-\beta}=\|T_{-\beta}(\mathbf{y})\|_{\infty,-\beta}\leq\tilde{C}\|y\|_{1,-\beta}=\tilde{C}\mu_{n}^{-\beta}\nu_{n}\|v\|,$
for every $m<n$. This easily implies that
$\|\mathcal{A}(m,n)v\|\leq\tilde{C}\left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}\|v\|,$
for every $m<n$ as claimed. ∎
For $m,n\in\mathbb{Z}$, set
$\mathcal{G}(m,n):=\begin{cases}\mathcal{A}(m,n)P_{n}&m\geq n;\\ -\mathcal{A}(m,n)(\text{Id}-P_{n})&m<n.\end{cases}$
Then, by Lemmas 4.7, 4.8 and 4.9 there exists $C^{\prime}>0$ such that
$\|\mathcal{G}(m,n)\|\leq C^{\prime}\begin{cases}\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}^{2}&m\geq n;\\ \left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}^{2}&m<n.\end{cases}$ (29)
Lemma 4.10.
There exists $C^{\prime\prime}>0$ such that
$\|\mathcal{G}(m,n)\|\leq C^{\prime\prime}\begin{cases}\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}&m\geq n;\\ \left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}&m<n.\end{cases}$
Proof of the Lemma 4.10.
Take $n\in{\mathbb{Z}}$ and $v\in X$. We define a sequence $\mathbf{y}=(y_{m})_{m\in\mathbb{Z}}$ by $y_{n}=v$ and $y_{m}=0$ for $m\neq n$. Then, $\mathbf{y}\in\ell_{\beta}^{1}\cap\ell_{-\beta}^{1}$. Set
$x_{m}:=\mathcal{G}(m,n)v,\quad m\in\mathbb{Z}.$
It is easy to verify that (23) holds. We claim that $\mathbf{x}=(x_{m})_{m\in\mathbb{Z}}\in\ell_{\beta}^{\infty}$. Indeed, by (29), for $m\geq n$ we have that
$\|x_{m}\|=\|\mathcal{G}(m,n)v\|\leq C^{\prime}\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n}^{2}\|v\|,$
which implies that
$\sup_{m\geq n}(\mu_{m}^{\beta}\|x_{m}\|)<+\infty.$
Similarly, for $m<n$ we have that
$\|x_{m}\|=\|\mathcal{G}(m,n)v\|\leq C^{\prime}\left(\frac{\mu_{n}}{\mu_{m}}\right)^{-\beta}\nu_{n}^{2}\|v\|,$
which implies that
$\mu_{m}^{-\beta}\|x_{m}\|\leq C^{\prime}\mu_{n}^{-\beta}\nu_{n}^{2}\|v\|.$
Thus, since $\lim_{m\to-\infty}\mu_{m}=0$ and, in particular, $\mu_{m}^{\beta}\leq\mu_{m}^{-\beta}$ for $m$ sufficiently small, the previous inequality implies that
$\sup_{m<n}(\mu_{m}^{\beta}\|x_{m}\|)<+\infty.$
Combining these observations we conclude that $\mathbf{x}=(x_{m})_{m\in\mathbb{Z}}\in\ell_{\beta}^{\infty}$ and $T_{\beta}(\mathbf{y})=\mathbf{x}$, where $T_{\beta}$ is as in Lemma 4.4. Therefore, for $m\geq n$ we have that
$\mu_{m}^{\beta}\|\mathcal{G}(m,n)v\|=\mu_{m}^{\beta}\|x_{m}\|\leq\|\mathbf{x}\|_{\infty,\beta}\leq\|T_{\beta}\|\cdot\|\mathbf{y}\|_{1,\beta}=\|T_{\beta}\|\mu_{n}^{\beta}\nu_{n}\|v\|,$
and thus
$\|\mathcal{G}(m,n)\|\leq\|T_{\beta}\|\left(\frac{\mu_{m}}{\mu_{n}}\right)^{-\beta}\nu_{n},\quad m\geq n.$
Similarly, one can treat the case when $m<n$. ∎
Now the proof of Theorem 4.2 can be easily completed by combining the previous auxiliary results. ∎
4.2. Persistence of $(\mu,\nu)$-dichotomy
We now present a version of Theorem 3.12 in the case of two-sided dynamics.
Theorem 4.11.
Let $(\gamma_{n})_{n\in{\mathbb{Z}}}$ be a sequence of numbers with $\gamma_{n}>0$ for every $n\in{\mathbb{Z}}$ such that
$\sum_{n\in{\mathbb{Z}}}\gamma_{n}<+\infty.$
Suppose that (2) admits a $(\mu,\nu)$-dichotomy with $\lambda>0$ and condition (22) is satisfied and take $\beta\in(0,\lambda-\varepsilon)$. Moreover, let $(B_{n})_{n\in{\mathbb{Z}}}$ be a sequence of operators in $\mathcal{B}(X)$ with the property that there exists $c>0$ such that
$\|B_{n}\|\leq\frac{c\gamma_{n}\mu_{n}^{\beta}}{\nu_{n+1}\mu_{n+1}^{\beta}},\quad\text{for $n\in\mathbb{Z}$.}$ (30)
Then, if $c$ is small enough we have that the nonautonomous difference equation
$x_{n+1}=(A_{n}+B_{n})x_{n},\quad n\in{\mathbb{Z}}$
also admits a $(\mu,\nu)$-dichotomy.
Proof.
The proof of this result is similar to the proof of Theorem 3.12 and, therefore, we refrain from writing it. ∎
Remark 4.12.
A comment similar to Remark 3.13 also applies to condition (30).
Acknowledgements. We would like to thank the referees for their useful comments that helped us to improve our paper and to C. Silva for useful discussions related to [44]. D.D is also grateful to W. Zhang and L. Zhou for ideas related to [20, 21] which inspired results of the present paper. L. Backes was partially supported by a CNPq-Brazil PQ fellowship under Grant No. 307633/2021-7. D. Dragičević was supported in part by University of Rijeka under the project uniri-iskusni-prirod-23-98 3046.
Statements and Declarations
Competing Interests: no potential conflict of interest was reported by the authors.